OpenStudy (anonymous):

Completely factor the expression 98x^2+84xy+18y^2

OpenStudy (radar):

2(7x+3y)^2

OpenStudy (radar):

2(7x+3y)(7x+3y) another way of saying the same thing.

OpenStudy (anonymous):

o ok thank you

OpenStudy (radar):

ur welcome

OpenStudy (anonymous):

Just more explanation: = 2 ( 49x² + 42xy + 9y²) = 2 [ ( 7x)² +42xy + (3y)² ] = 2 [( 7x)+ (3y) ]²

OpenStudy (radar):

Thanks for showing more detailed steps than I. Hopefully dawn354215 sees the relationships.

OpenStudy (anonymous):

I'm still learning, people in OS are way intelleligent, rapidly and sophisticated --> I'm addicted to the enviroment :)